# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ VI Cavity, rates, and master equation
We now promote the single-pass excitation probabilities into rates by (i) specifying how the cavity selects one outgoing Schwarzschild mode and how much of it escapes to infinity through the exterior potential (greybody factor), and (ii) dividing by the relevant interaction time associated with the atom’s transit across the mode region. The Proca field introduces polarization-dependent greybody transmission: the axial channel is single, while the polar sector yields two effective channels after local diagonalization. Near the horizon all channels share the same plane-wave phase responsible for the thermal factor; the greybody matrices merely supply smooth, frequency-dependent multipliers in the prefactor and impose the mass threshold $\nu\geq m_{V}$. These ingredients feed directly into the emission/absorption rates used in the master equation.
@@ VI.1 Mode selection and greybody factors
For each $\ell$ the master radial equations of Section III define an elastic scattering problem on the line of $r_{*}$. Let $\Psi^{\text{in}}_{\ell}$ denote a unit-flux wave incident from the near-horizon side. The Wronskian constancy fixes the scattering matrix and the transmission coefficient $\mathcal{T}_{\ell}(\nu)\in[0,1]$. For the Proca field, the axial sector provides one channel with transmission $\mathcal{T}_{\ell}^{\text{(ax)}}(\nu)$ for $\ell\geq 1$. On the other hand, the polar sector yields two effective channels after local diagonalization with transmissions $\mathcal{T}_{\ell}^{\text{(pol)}\pm}(\nu)$ (for $\ell\geq 0$, noting that $\ell=0$ is polar only).
The single-mode escape factor multiplying probabilities is then
$\Xi_{\ell}(\nu)=\begin{cases}\mathcal{T}_{\ell}^{\text{(ax)}}(\nu),&\text{axial channel},\\ \mathcal{T}_{\ell}^{\text{(pol)}+}(\nu)+\mathcal{T}_{\ell}^{\text{(pol)}-}(\nu),&\text{polar sector}.\end{cases}$ (81)
Near the horizon, the axial/polar distinction is immaterial for the thermal factor; it re-enters only through $\Xi_{\ell}(\nu)$ as a smooth function of $\nu$.
Two universal limits will be repeatedly used:
$\Xi_{\ell}(\nu)\rightarrow[\nu\to m_{V}^{+}]{}0,\qquad\Xi_{\ell}(\nu)\rightarrow[\nu\gg\max{m_{V},\ell/r_{g}}]{}1.$ (82)
@@
$P_{\text{exc,esc}}^{\bullet}(\nu,\ell)=\frac{4\pi\,\nu}{\omega^{2}}\,\frac{\mathcal{K}_{\bullet}}{e^{4\pi\nu}-1}\,\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\underbrace{\Xi_{\ell}(\nu)\,\overline{\mathcal{P}}_{\bullet}(\nu,\ell)}_{\text{smooth prefactor}},$ (83)
where $\mathcal{K}_{\bullet}$ is the detector coupling factor and $\overline{\mathcal{P}}_{\bullet}$ is the polarization/geometry weight from Section IV. The factor $\Xi_{\ell}(\nu)$ will remain attached when we convert probabilities to rates and then to the master equation, ensuring that only the fraction of quanta that traverse the exterior potential contribute to the outgoing flux at $\mathscr{I}^{+}$. Noting explicitly, $\Xi_{\ell}(\nu)\sim k_{\infty}^{2\ell+1}$ near threshold generically, so $\Xi_{\ell}/k_{\infty}\sim k_{\infty}^{2\ell}$ is finite/vanishing (and never divergent).
@@ VI.2 Emission and absorption rates
We now convert the single-pass escaping probabilities in (83) into mode rates by multiplying by the atomic throughput across the cavity. Let $R_{\!a}$ denote the number of atoms (prepared in $|g\rangle)$ entering the cavity per unit Killing time $t$ (dimensionless as in Section II). For a cavity-selected outgoing channel $(\nu,\ell)$ the emission rate is
$\Gamma_{e}^{\bullet}(\nu,\ell)=R_{\!a}\;P_{\text{exc,esc}}^{\bullet}(\nu,\ell)=R_{\!a}\,\frac{4\pi\,\nu}{\omega^{2}}\,\frac{\mathcal{K}_{\bullet}}{e^{4\pi\nu}-1}\,\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\Xi_{\ell}(\nu)\,\overline{\mathcal{P}}_{\bullet}(\nu,\ell)\;,$ (84)
with $k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}}$, $\mathcal{K}_{\text{mono}}=g_{V}^{2}|m_{eg}|^{2}$, $\mathcal{K}_{\text{dip}}=g_{d}^{2}|d_{eg}|^{2}$, and $\overline{\mathcal{P}}_{\bullet}$ the smooth polarization/geometry factor defined in Section IV. The absorption rate follows from the universal detailed balance established in Section V,
$\Gamma_{a}^{\bullet}(\nu,\ell)=e^{\,4\pi\nu}\,\Gamma_{e}^{\bullet}(\nu,\ell);\,$ (85)
so the ratio $\Gamma_{e}^{\bullet}/\Gamma_{a}^{\bullet}=e^{-4\pi\nu}$ is independent of detector realization, mass, or greybody weights Kubo:1957mj; Martin:1959jp; breuer2002theory. Note that $R_{a}$ affects absolute fluxes and relaxation speeds, but cancels from $\langle n\rangle_{\mathrm{ss}}$.
@@
$\kappa_{\bullet}\;\equiv\;R_{\!a}\,\mathcal{K}_{\bullet},$ (86)
@@
[…]
@@
$\Gamma_{e,a}^{\bullet}(\nu)=\sum_{\ell}\Gamma_{e,a}^{\bullet}(\nu,\ell),$ (88)
bearing in mind that axial/polar channels contribute through their respective $\Xi_{\ell}(\nu)$ and (implicitly) through $\overline{\mathcal{P}}_{\bullet}(\nu,\ell)$. In the massless limit $m_{V}\to 0$ one has $k_{\infty}\to\nu$ and the threshold factor becomes trivial, recovering the familiar $\propto\nu/\omega^{2}$ scaling of the scalar-like response (up to the polarization prefactor). These $\Gamma_{e,a}$ are the inputs for the mode-occupancy master equation in the next subsection.
@@ VI.3 Master equation and steady state
For the single, cavity-selected outgoing mode of frequency $\nu$, we model the coarse-grained field dynamics with a birth-death master equation for the diagonal elements $\rho_{nn}(t)\equiv\langle n|\rho(t)|n\rangle$ in the Fock basis. With emission and absorption rates $\Gamma_{e}^{\bullet}(\nu,\ell)$ and $\Gamma_{a}^{\bullet}(\nu,\ell)$ from (87), the population dynamics reads
$\frac{d\rho_{nn}}{dt}=-\,\Gamma_{e}\left[(n+1)\rho_{nn}-n\,\rho_{n-1,n-1}\right]-\,\Gamma_{a}\left[n\,\rho_{nn}-(n+1)\rho_{n+1,n+1}\right],$ (89)
where, to declutter notation, we have fixed $(\nu,\ell)$ and the detector label $\bullet$ and written $\Gamma_{e,a}\equiv\Gamma_{e,a}^{\bullet}(\nu,\ell)$. This equation is equivalent to the standard Lindblad form with jump operators $a^{\dagger}$ and $a$ Lindblad:1975ef,
$\dot{\rho}\;=\;\Gamma_{e}\!\left(a^{\dagger}\rho a-\frac{1}{2}{aa^{\dagger}\,\rho}\right)+\Gamma_{a}\!\left(a\rho a^{\dagger}-\frac{1}{2}{a^{\dagger}a\,\rho}\right),$ (90)
@@
The detailed-balance relation proved in (78)-(79) implies $\Gamma_{e}/\Gamma_{a}=e^{-4\pi\nu}$. The steady state $\dot{\rho}_{nn}=0$ of (89) is therefore geometric:
$\frac{\rho_{n+1,n+1}^{(\text{ss})}}{\rho_{nn}^{(\text{ss})}}=\frac{\Gamma_{e}}{\Gamma_{a}}=e^{-4\pi\nu}\;\Longrightarrow\;\rho_{nn}^{(\text{ss})}=(1-e^{-4\pi\nu})\,e^{-4\pi\nu\,n}.$ (91)
The mean mode occupancy follows either from (89) or directly from (91):
$\frac{d\langle n\rangle}{dt}=-(\Gamma_{a}-\Gamma_{e})\,\langle n\rangle+\Gamma_{e},\qquad\;\langle n\rangle_{\text{ss}}=\frac{\Gamma_{e}}{\Gamma_{a}-\Gamma_{e}}=\frac{1}{e^{4\pi\nu}-1}\;.$ (92)
@@
$\langle n(t)\rangle-\langle n\rangle_{\text{ss}}=\left(\langle n(0)\rangle-\langle n\rangle_{\text{ss}}\right)\,e^{-(\Gamma_{a}-\Gamma_{e})t}.$ (93)
When several $\ell$-channels are admitted, the total master equation is a direct sum of copies of (89), one per channel, with $\Gamma_{e,a}(\nu,\ell)$ as in (87). Each channel relaxes to the same geometric law (91) with the same $\langle n\rangle_{\text{ss}}$ in (92), while differing only in relaxation rate and overall flux set by $\Xi_{\ell}(\nu)$, $k_{\infty}(\nu)$, and the detector-dependent prefactor carried in $\Gamma_{e,a}$.
@@ VII.1 Entropy flux from the master equation
$\frac{d\rho_{nn}}{dt}=-\,\Gamma_{e}\left[(n+1)\rho_{nn}-n\,\rho_{n-1,n-1}\right]-\,\Gamma_{a}\left[n\,\rho_{nn}-(n+1)\rho_{n+1,n+1}\right],$ (94)
@@
$\frac{d\langle n\rangle}{dt}=-(\Gamma_{a}-\Gamma_{e})\,\langle n\rangle+\Gamma_{e}.$ (95)
@@
$\dot{\bar{n}}_{\nu}\;\equiv\;\frac{d\langle n\rangle}{dt}.$ (96)
@@
$\frac{dS}{dt}=-k_{B}\sum_{n}\dot{\rho}_{nn}\,\ln\rho_{nn}.$ (97)
Following the standard near-stationary replacement, we evaluate the logarithm at the instantaneous steady state corresponding to the same rates in (94). The steady state solves $\Gamma_{e}/\Gamma_{a}=e^{-4\pi\nu}$ with
$\rho_{nn}^{(\mathrm{ss})}=(1-e^{-4\pi\nu})\,e^{-4\pi\nu\,n}.$ (98)
Using $\ln\rho_{nn}^{(\mathrm{ss})}=\ln(1-e^{-4\pi\nu})-4\pi\nu\,n$ in (97), the constant term drops by probability conservation $\sum_{n}\dot{\rho}_{nn}=0$. We obtain
$\frac{dS}{dt}=4\pi k_{B}\nu\sum_{n}n\,\dot{\rho}_{nn}=4\pi k_{B}\nu\,\frac{d\langle n\rangle}{dt}.$ (99)
@@
$\frac{dS_{p}}{dt}\;=\;4\pi k_{B}\,\nu\;\dot{\bar{n}}_{\nu}\;,$ (100)
where $\dot{\bar{n}}_{\nu}$ is determined by (95) with the Proca-specific rates from Section VI. The coefficient $4\pi\nu$ is universal: it depends only on the near-horizon mapping that fixes detailed balance, whereas the detector model, polarization content, greybody transmission, and the mass threshold $\nu\geq m_{V}$ enter exclusively through $\dot{\bar{n}}_{\nu}$. When several $\ell$-channels are admitted simultaneously, the total entropy flux is the sum of (100) over the admitted channels, each with its own $\dot{\bar{n}}_{\nu}(\ell)$.

### v2

@@ VI Cavity, rates, and master equation
We now promote the single-pass excitation probabilities into rates by (i) specifying how the cavity selects one outgoing Schwarzschild mode and how much of it escapes to infinity through the exterior potential (greybody factor), and (ii) dividing by the relevant interaction time associated with the atom’s transit across the mode region. The Proca field introduces polarization-dependent greybody transmission: the axial channel is single, while the polar sector yields two effective channels after local diagonalization. Near the horizon all channels share the same plane wave phase responsible for the thermal factor; the greybody matrices merely supply smooth, frequency-dependent multipliers in the prefactor and impose the mass threshold $\nu\geq m_{V}$. These ingredients feed directly into the emission/absorption rates used in the master equation.
@@ VI.1 Mode selection and greybody factors
For each $\ell$ the master radial equations of Sec. III define an elastic scattering problem on the line of $r_{*}$. Let $\Psi^{\text{in}}_{\ell}$ denote a unit-flux wave incident from the near-horizon side. The Wronskian constancy fixes the scattering matrix and the transmission coefficient $\mathcal{T}_{\ell}(\nu)\in[0,1]$. For the Proca field, the axial sector provides one channel with transmission $\mathcal{T}_{\ell}^{\text{(ax)}}(\nu)$ for $\ell\geq 1$. On the other hand, the polar sector yields two effective channels after local diagonalization with transmissions $\mathcal{T}_{\ell}^{\text{(pol)}\pm}(\nu)$ (for $\ell\geq 0$, noting that $\ell=0$ is polar only).
At this stage it is important to distinguish inclusion from explicit evaluation. The present paper does not solve the Proca transmission problem numerically for each $(\ell,\nu)$; rather, it formulates the escaping HBAR flux in terms of the exact transmission coefficients associated with the Schwarzschild radial scattering problem. This is sufficient for the purposes of the present analysis because the universal thermal factor was already fixed in Sec. V by the near-horizon worldline integral, whereas the exterior barrier enters only through the smooth escape factor defined below. The single-mode escape factor multiplying probabilities is then
$\Xi_{\ell}(\nu)=\begin{cases}\mathcal{T}_{\ell}^{\text{(ax)}}(\nu),&\text{axial channel},\\ \mathcal{T}_{\ell}^{\text{(pol)}+}(\nu)+\mathcal{T}_{\ell}^{\text{(pol)}-}(\nu),&\text{polar sector}.\end{cases}$ (VI.2)
Near the horizon, the axial/polar distinction is immaterial for the thermal factor; it reenters only through $\Xi_{\ell}(\nu)$ as a smooth function of $\nu$. Accordingly, the present work establishes how polarization-resolved greybody transmission modifies the HBAR prefactor and escaping flux in full generality, while leaving a dedicated numerical evaluation of the axial/polar transmission functions for future study.
Two universal limits will be repeatedly used
$\Xi_{\ell}(\nu)\rightarrow[\nu\to m_{V}^{+}]{}0,\qquad\Xi_{\ell}(\nu)\rightarrow[\nu\gg\max{m_{V},\ell/r_{g}}]{}1.$ (VI.3)
@@
$P_{\text{exc,esc}}^{\bullet}(\nu,\ell)=\frac{4\pi\,\nu}{\omega^{2}}\,\frac{\mathcal{K}_{\bullet}}{e^{4\pi\nu}-1}\,\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\underbrace{\Xi_{\ell}(\nu)\,\overline{\mathcal{P}}_{\bullet}(\nu,\ell)}_{\text{smooth prefactor}},$ (VI.4)
where $\mathcal{K}_{\bullet}$ is the detector coupling factor and $\overline{\mathcal{P}}_{\bullet}$ is the polarization/geometry weight from Sec. IV. The factor $\Xi_{\ell}(\nu)$ will remain attached when we convert probabilities to rates and then to the master equation, ensuring that only the fraction of quanta that traverse the exterior potential contribute to the outgoing flux at $\mathscr{I}^{+}$. Noting explicitly, $\Xi_{\ell}(\nu)\sim k_{\infty}^{2\ell+1}$ near threshold generically, so $\Xi_{\ell}/k_{\infty}\sim k_{\infty}^{2\ell}$ is finite/vanishing (and never divergent).
@@ VI.2 Emission and absorption rates
We now convert the single-pass escaping probabilities in (VI.4) into mode rates by multiplying by the atomic throughput across the cavity. Let $R_{\!a}$ denote the number of atoms (prepared in $|g\rangle)$ entering the cavity per unit Killing time $t$ (dimensionless as in Sec. II). For a cavity-selected outgoing channel $(\nu,\ell)$ the emission rate is
$\Gamma_{e}^{\bullet}(\nu,\ell)=R_{\!a}\;P_{\text{exc,esc}}^{\bullet}(\nu,\ell)$
$=R_{\!a}\,\frac{4\pi\,\nu}{\omega^{2}}\,\frac{\mathcal{K}_{\bullet}}{e^{4\pi\nu}-1}\,\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\Xi_{\ell}(\nu)\,\overline{\mathcal{P}}_{\bullet}(\nu,\ell)\;,$ (VI.5)
with $k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}}$, $\mathcal{K}_{\text{mono}}=g_{V}^{2}|m_{eg}|^{2}$, $\mathcal{K}_{\text{dip}}=g_{d}^{2}|d_{eg}|^{2}$, and $\overline{\mathcal{P}}_{\bullet}$ the smooth polarization/geometry factor defined in Sec. IV. The absorption rate follows from the universal detailed balance established in Sec. V,
$\Gamma_{a}^{\bullet}(\nu,\ell)=e^{\,4\pi\nu}\,\Gamma_{e}^{\bullet}(\nu,\ell);\,$ (VI.6)
so the ratio $\Gamma_{e}^{\bullet}/\Gamma_{a}^{\bullet}=e^{-4\pi\nu}$ is independent of detector realization, mass, or greybody weights [34, 36, 11]. Note that $R_{a}$ affects absolute fluxes and relaxation speeds, but cancels from $\langle n\rangle_{\mathrm{ss}}$.
@@
$\kappa_{\bullet}\;\equiv\;R_{\!a}\,\mathcal{K}_{\bullet},$ (VI.7)
@@
$\Gamma_{e}^{\bullet}(\nu,\ell)$ $=\frac{4\pi\,\nu}{\omega^{2}}\,\frac{\kappa_{\bullet}}{e^{4\pi\nu}-1}\,\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\Xi_{\ell}(\nu)\,\overline{\mathcal{P}}_{\bullet}(\nu,\ell),$
$\Gamma_{a}^{\bullet}(\nu,\ell)$ $=e^{\,4\pi\nu}\,\Gamma_{e}^{\bullet}(\nu,\ell).$ (VI.8)
@@
$\Gamma_{e,a}^{\bullet}(\nu)=\sum_{\ell}\Gamma_{e,a}^{\bullet}(\nu,\ell),$ (VI.9)
bearing in mind that axial/polar channels contribute through their respective $\Xi_{\ell}(\nu)$ and (implicitly) through $\overline{\mathcal{P}}_{\bullet}(\nu,\ell)$. In the massless limit $m_{V}\to 0$ one has $k_{\infty}\to\nu$ and the threshold factor becomes trivial, recovering the familiar $\propto\nu/\omega^{2}$ scaling of the scalarlike response (up to the polarization prefactor). These $\Gamma_{e,a}$ are the inputs for the mode-occupancy master equation in the next subsection.
@@ VI.3 Master equation and steady state
For the single, cavity-selected outgoing mode of frequency $\nu$, we model the coarse-grained field dynamics with a birth-death master equation for the diagonal elements $\rho_{nn}(t)\equiv\langle n|\rho(t)|n\rangle$ in the Fock basis. With emission and absorption rates $\Gamma_{e}^{\bullet}(\nu,\ell)$ and $\Gamma_{a}^{\bullet}(\nu,\ell)$ from (VI.8), the population dynamics reads
$\frac{d\rho_{nn}}{dt}$ $=-\,\Gamma_{e}\left[(n+1)\rho_{nn}-n\,\rho_{n-1,n-1}\right]$
$-\,\Gamma_{a}\left[n\,\rho_{nn}-(n+1)\rho_{n+1,n+1}\right],$ (VI.10)
where, to declutter notation, we have fixed $(\nu,\ell)$ and the detector label $\bullet$ and written $\Gamma_{e,a}\equiv\Gamma_{e,a}^{\bullet}(\nu,\ell)$. This equation is equivalent to the standard Lindblad form with jump operators $a^{\dagger}$ and $a$ [35],
$\dot{\rho}\;=\;\Gamma_{e}\!\left(a^{\dagger}\rho a-\frac{1}{2}{aa^{\dagger}\,\rho}\right)+\Gamma_{a}\!\left(a\rho a^{\dagger}-\frac{1}{2}{a^{\dagger}a\,\rho}\right),$ (VI.11)
@@
The detailed-balance relation proved in (V.19)–(V.20) implies $\Gamma_{e}/\Gamma_{a}=e^{-4\pi\nu}$. The steady state $\dot{\rho}_{nn}=0$ of (VI.10) is therefore geometric:
$\frac{\rho_{n+1,n+1}^{(\text{ss})}}{\rho_{nn}^{(\text{ss})}}=\frac{\Gamma_{e}}{\Gamma_{a}}=e^{-4\pi\nu}\;\Longrightarrow\;\rho_{nn}^{(\text{ss})}=(1-e^{-4\pi\nu})\,e^{-4\pi\nu\,n}.$ (VI.12)
The mean mode occupancy follows either from (VI.10) or directly from (VI.12):
$\frac{d\langle n\rangle}{dt}$ $=-(\Gamma_{a}-\Gamma_{e})\,\langle n\rangle+\Gamma_{e},$
$\;\langle n\rangle_{\text{ss}}$ $=\frac{\Gamma_{e}}{\Gamma_{a}-\Gamma_{e}}=\frac{1}{e^{4\pi\nu}-1}\;.$ (VI.13)
@@
$\langle n(t)\rangle-\langle n\rangle_{\text{ss}}=\left(\langle n(0)\rangle-\langle n\rangle_{\text{ss}}\right)\,e^{-(\Gamma_{a}-\Gamma_{e})t}.$ (VI.14)
When several $\ell$-channels are admitted, the total master equation is a direct sum of copies of (VI.10), one per channel, with $\Gamma_{e,a}(\nu,\ell)$ as in (VI.8). Each channel relaxes to the same geometric law (VI.12) with the same $\langle n\rangle_{\text{ss}}$ in (VI.13), while differing only in relaxation rate and overall flux set by $\Xi_{\ell}(\nu)$, $k_{\infty}(\nu)$, and the detector-dependent prefactor carried in $\Gamma_{e,a}$.
@@ VII.1 Entropy flux from the master equation
$\frac{d\rho_{nn}}{dt}$ $=-\,\Gamma_{e}\left[(n+1)\rho_{nn}-n\,\rho_{n-1,n-1}\right]$
$-\,\Gamma_{a}\left[n\,\rho_{nn}-(n+1)\rho_{n+1,n+1}\right],$ (VII.1)
@@
$\frac{d\langle n\rangle}{dt}=-(\Gamma_{a}-\Gamma_{e})\,\langle n\rangle+\Gamma_{e}.$ (VII.2)
@@
$\dot{\bar{n}}_{\nu}\;\equiv\;\frac{d\langle n\rangle}{dt}.$ (VII.3)
@@
$\frac{dS}{dt}=-k_{B}\sum_{n}\dot{\rho}_{nn}\,\ln\rho_{nn}.$ (VII.4)
Following the standard near-stationary replacement, we evaluate the logarithm at the instantaneous steady state corresponding to the same rates in (VII.1). The steady state solves $\Gamma_{e}/\Gamma_{a}=e^{-4\pi\nu}$ with
$\rho_{nn}^{(\mathrm{ss})}=(1-e^{-4\pi\nu})\,e^{-4\pi\nu\,n}.$ (VII.5)
Using $\ln\rho_{nn}^{(\mathrm{ss})}=\ln(1-e^{-4\pi\nu})-4\pi\nu\,n$ in (VII.4), the constant term drops by probability conservation $\sum_{n}\dot{\rho}_{nn}=0$. We obtain
$\frac{dS}{dt}=4\pi k_{B}\nu\sum_{n}n\,\dot{\rho}_{nn}=4\pi k_{B}\nu\,\frac{d\langle n\rangle}{dt}.$ (VII.6)
@@
$\frac{dS_{p}}{dt}\;=\;4\pi k_{B}\,\nu\;\dot{\bar{n}}_{\nu}\;,$ (VII.7)
where $\dot{\bar{n}}_{\nu}$ is determined by (VII.2) with the Proca-specific rates from Sec. VI. The coefficient $4\pi\nu$ is universal: it depends only on the near-horizon mapping that fixes detailed balance, whereas the detector model, polarization content, greybody transmission, and the mass threshold $\nu\geq m_{V}$ enter exclusively through $\dot{\bar{n}}_{\nu}$. When several $\ell$-channels are admitted simultaneously, the total entropy flux is the sum of (VII.7) over the admitted channels, each with its own $\dot{\bar{n}}_{\nu}(\ell)$.